### v3

# Partial sampling of a random spanning tree

Abstract
We present a decomposition of the distribution of the subtree connecting several vertices for spanning trees sampled in a complete graph and a short self-contained derivation of its asymptotic behavior.
00footnotetext: Key words and phrases: Trees, complete graphs00footnotetext: AMS 2020 subject classification: 60C05, 60J10
1 Introduction
The purpose of this note is to address the following simple question: Given a random spanning tree $\Upsilon$ of the complete graph with $n$ vertices, and a fixed subset $L$ of $d<n$ specific labeled vertices, how are these $d$ vertices connected within the tree when $n$ is large? More precisely, considering the topological type of the smallest subtree $\Upsilon_{L}$ of the random (uniformly sampled) spanning tree connecting them and the graph distances between all leaves and nodes of this subtree, what is their asymptotic distribution as $n$ increases to infinity?
The theorem given below provides an answer in a slightly more general situation: One of the specific vertices denoted by $\Delta$ is chosen to be the root of $\Upsilon$. The weight of a tree is assumed to be proportional to $\kappa^{p}$, $\kappa$ being a positive parameter and $p$ the degree of $\Delta$ in the spanning tree. For $\kappa\neq 1$, the distribution we get on spanning tree is not entirely uniform.
However it turns out that the choice of $\kappa$ is irrelevant to answer the question: We show that asymptotically, the topology of $\Upsilon_{L}$ is uniformly chosen among binary trees whose leaves are $L$, and that the graph distances between all leaves and nodes, normalized by $\sqrt{n}$ , converge towards an explicit distribution. As $d$ increases, these distributions (necessarily) form a consistent system.
The question of constructing $\Upsilon_{L}$ can be raised of course on any graph and we show in section 2 how Wilson’s algorithm allows to solve it in full generality. Once objects have been rigorously defined and notations established, the convergence proof given in section 3 has the merit to be short and simple.
As far as we know, this result did not appear as such in the literature. However, it is very close of a description of the Continuous Random Tree given by Aldous in [3]. The consistent system of asymptotic distributions we get can certainly be interpreted in terms of sampling of the CRT. The point here is to show their relevance at the discrete elementary level. More details and references are given at the end of the last section.
2 Definitions, first properties and statement
Consider a graph $\mathcal{G}$ equipped with conductances and a non-vanishing killing function defined on the set of vertices, here denoted by $X$. These data allow to define a Green matrix $G$ indexed by pairs of vertices. Adding a cemetery point $\Delta$, we can consider the value of the killing function at any vertex $x$ as a conductance between $x$ and $\Delta$. Then a spanning forest of rooted trees on $\mathcal{G}$ can be identified to a spanning tree on the extended graph $\mathcal{G}_{\Delta}$, by connecting $\Delta$ to the roots. An extension of Cayley’s theorem (cf. for example [7], section 8-2 in [8]) shows that if we define the weight of such a spanning tree to be the product of the conductances of its edges, the sum of these weights is the determinant of $G$. It provides naturally a probability $\mathbb{P}$ on spanning trees rooted in $\Delta$ and consequently a probability on spanning forests.
This applies in particular to the spanning forests of the complete graph $K_{n}$ with vertices $\{1,2,...,n\}$ endowed with unit conductances and a constant killing function $\kappa$. The corresponding probability on rooted trees will be denoted by $\mathbb{P}^{(n),\kappa}$. If $\kappa=1$, the spanning tree rooted in $\Delta$ can be identified with a uniform spanning tree on $K_{n+1}$. Local limits for these objects have been determined in [4], extending a result of Grimmett [5].
The expression of the Green matrix $G^{(n),\kappa}$ is:
$\frac{1}{\kappa(n+\kappa)}(\kappa I+J)$
where $J$ denotes the $(n,n)$ matrix with all entries equal to $1$. It is easy to check that for any $(m,m)$ square matrix $M$ with diagonal entries equal to $a+b$ and off diagonal entries equal to $b$, $\det(M)=a^{m-1}(a+mb)$. Hence we get the following identity:
$\det(G^{(n),\kappa}=\frac{1}{\kappa(n+\kappa)^{n-1}}$ (1)
Moreover, for $0<d<n$,
$\det(G^{(n),\kappa}_{i,j},\,1\leq i,j\leq d)=\frac{d+\kappa}{\kappa(n+\kappa)^{d}}$ (2)
Coming back to the general case, a simple way to sample a spanning tree $\Upsilon$ under the probability $\mathbb{P}$ is to perform Wilson’s algorithm (cf. [12]) which is based on loop erasure. The probability of a loop-erased path from a given vertex $\xi_{0}$ to $\Delta$ is the product of its conductances normalized by the determinant of the restriction of the Green matrix to the vertices of $\xi$. Wilson’s algorithm starts by choosing any order on the vertices and constructing a loop erased path from the first vertex to $\Delta$, then at each step it constructs a loop erased path from the first unused vertex to the tree of used vertices (including $\Delta$) until all vertices have been used.
We are interested in the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ connecting a set $L$ of $l$ vertices and $\Delta$. For that, we can run Wilson algorithm after choosing an order in which these vertices are the first $l$ vertices. Given a tree $Y$ imbedded in the graph and rooted in $\Delta$ containing $L$ and whose set of leaves is included in $L$ , the spanning tree $\Upsilon$ produced by the algorithm will contain it iff before the $l$-th step of the algorithm the tree of used vertices is exactly $Y$. Then we will have $\Upsilon_{L}=Y$.
It is easy to check from the proof of Wilson’s algorithm given in [8] that for any tree $Y$ rooted in $\Delta$, the probability $\mathbb{P}(Y=\Upsilon_{L})$ is given by the product of the conductances of the edges of $Y$ normalized by the determinant of the restriction of the Green matrix to the vertices of $Y$.
Consequently, in the case of the complete graph with $n$ vertices $\{1,...,n\}$ endowed with unit conductances and a constant killing factor $\kappa$ if $Y$ has $d$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
Lemma 2.1
$\mathbb{P}^{(n),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(d+\kappa)}{(n+\kappa)^{d}}$
A spanning tree defines a map $p$ from $X$ into $X\cup\{\Delta\}$ fixing $\Delta$ and such that for any vertex $x$ in $X$ and positive integer $m$, $p^{m}(x)\neq x$. If $p(y)=x$, we say that $y$ is a child of $x$. If $p^{m}(y)=x$ for some non-negative (positive) integer $m$, we say that $y$ is above (strictly above) $x$ and that $x$ is below (strictly below) $y$. A node is by definition a vertex with at least two children. These notions are extended to any tree imbedded in the graph and rooted in $\Delta$ when we replace $X$ by the set of tree vertices.
Let us come back to the case of the complete graph $K_{n}$.
We say that a tree imbedded in the complete graph and rooted in $\Delta$ is a $L$-tree if it contains $L$ and his set of leaves, denoted by $L^{*}$, is contained in $L$. $\Upsilon_{L}$ is always a $L$-tree. We denote by $N$ the set of nodes of $\Upsilon_{L}$. There is a mapping $j$ from $N\cup L$ into parts of $L$. It maps a vertex $x$ to the set of elements of $L$ which are above it in the tree. In particular, it sends a leaf $i$ into $\{i\}$. We also set $j(\Delta)=L$.
We say that a tree is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. The order on $L$ is used to induce an order on subsets of $L$ by listing the leaves in the set in increasing order and then use alphabetical order. For example we have $1<(1,3,7)<2<(2,3)$. Nodes and leaves of a reduced $L$-tree inherit of that order we will call the $L$-order.
In order to capture the way elements of $L$ are connected, in other words the geometry of the subtree connecting them, only the labels of $L$ do matter. This leads us to define an equivalence relation: we say that two $L$-trees are equivalent if they can be exchanged by a permutation of $\{l+1,...,n\}$. Note that equivalence preserves the $L$-order we defined on $N\cup L$. The (equivalence) class of the $L$-reduction of $\Upsilon_{L}$ denoted by $Q_{L}$ inherits a tree structure. Its inner vertices are the images of $N\cup L\setminus L^{*}$ by $j$ and its leaves singletons in $L^{*}$. The class of $\Upsilon_{L}$ is determined by $Q_{L}$ and the numbers of intermediate inner vertices between each vertex in $N\cup L$ and the first vertex in $N\cup L\setminus L^{*}\cup\{\Delta\}$ below it, hence by a map $u_{L}$ from the set of vertices of $Q_{L}$ into the set $\mathbb{N}$ of nonnegative integers. We can say that $(Q_{L},u_{L})$ is a $\mathbb{N}$-extension of $Q_{L}$.
We say that a $L$ tree is binary iff $L^{*}=L$, every internal vertex has exactly two children and the root only one. An easy induction shows that the set $\mathcal{B}_{l}$ of binary $L$-tree classes has cardinality $c_{l}=\prod_{i=1}^{l-1}(2i-1)=\frac{(2l-3)!}{2^{l-2}(l-2)!}$ and that these trees have l1l-1 internal vertices. Indeed, adding one leaf is done by choosing a vertex, add a node just below it and connect it to the new leaf. Consider now binary tree extensions: using the $L$-order on leaves and nodes, $u_{L}$ becomes a (2l1)(2l-1)-tuple of nonnegative integers $(u_{L}(i),1\leq i\leq 2l-1)$.
We can now formulate precisely our problem which is to show that as $n$ increases to infinity, the probability that the reduction of $\Upsilon_{L}$ is binary converges to one, and to determine the asymptotic behavior of the joint distribution of the pair $(Q_{L},u_{L})$.
The answer is given in the following:
Theorem 2.1
a) $\mathbb{P}^{(n),\kappa}(L^{*}=L)$ converges to 1 as $n\uparrow\infty$.
b) Given $\alpha\in\mathcal{B}_{l}$,
$\lim_{n\uparrow\infty}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha)=c_{l}^{-1}.$
c) Given $t\in\mathbb{R}_{+}^{2l-1}$,
$\lim_{n\uparrow\infty}n^{(2l-1)/2}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{n}])=(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}$
in which $t\sqrt{n}]$ denotes the 2l12l-1-tuple of integers $([\sqrt{n}t_{1}],...,[\sqrt{n}t_{2l-1}])$.
3 Proof
Note that a) follows directly from b) since there are $c_{l}$ binary trees with $l$ labeled leaves.
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{n}]+l-1$, there are $\prod_{i=l}^{\Sigma+l}(n-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{n}]$. From lemma 2.1, we get that
$\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{n}])=\frac{\Sigma+l+\kappa}{(n+\kappa)^{\Sigma+l}}\prod_{i=l}^{\Sigma+l}(n-i).$
Using Stirling’s approximation, it follows that
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}]))=\ln(\Sigma+l+\kappa)-(\Sigma+l)\ln(n+\kappa)+(n-l)\ln(n-l)-n+l+\frac{1}{2}\ln(2\pi(n-l))-(n-\Sigma-l)\ln(n-\Sigma-l)+n-\Sigma-l-\frac{1}{2}\ln(2\pi(n-\Sigma-l))+O(1/n)$
which, after decomposing the first term into $1/2\ln(n)+\ln(\sum t_{i})+\ln((\Sigma+l+\kappa)/(\sum t_{i}\sqrt{n}))$ and each term of the form $\ln(n+a)$ into $\ln(n)+\ln(1+a/n)$, by Taylor’s formula applied to the $\ln$ functions gives:
$\ln(\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}(i)=[t_{i}\sqrt{n}])=1/2\ln(n)+\ln(\sum t_{i})-l\ln(n)-\Sigma+n\ln(1-\frac{\Sigma-l}{n})+O(n^{-1/2})=\ln(\Sigma/\sqrt{n})-(2l-1)\ln(\sqrt{n})-\Sigma^{2}/2n+O(n^{-1/2}).$
The statement c) follows. One notes that if the $t_{i}$’s are bounded, the remainder term $O(1/n)$ in the first equation can be bounded uniformly by a constant multiple of $1/n$. Similarly, in the last equation, $O(n^{-1/2})$ can be uniformly bounded by a constant multiple of $1/\sqrt{n}$.
To prove b), we can first check that $\int_{\mathbb{R}_{+}^{2l-1}}(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}=c_{l}^{-1}$. Then the following Riemann sum argument shows that b) follows from c): the remark ending the proof of c) shows that given $k\in\mathbb{N}^{2l-1}$ such that $k_{i}<c\sqrt{n}$, $\lim_{n\uparrow\infty}n^{(2l-1)/2}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=k)=(\sum_{1}^{2l-1}k_{i}/\sqrt{n})e^{-(\sum_{1}^{2l-1}k_{i})^{2}/2n}$ uniformly in $k$.
Hence $\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha)>\lim_{n\uparrow\infty}\sum_{k\in\mathbb{N}^{2l-1},k_{i}<c\sqrt{n}}\mathbb{P}^{(n),\kappa}(Q_{L}=\alpha,u_{L}=k)=n^{-(2l-1)/2}\sum_{k\in\mathbb{N}^{2l-1},k_{i}<c\sqrt{n}}(\sum_{1}^{2l-1}k_{i}/\sqrt{n})e^{-(\sum_{1}^{2l-1}k_{i})^{2}/2n}+O(n^{-1/2})$˙It converges towards $\int_{[0,c]^{2l-1}}(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}\prod_{i}dt_{i}$ which can be made arbitrarily close from $c_{l}^{-1}$ by taking $c$ large enough. This concludes the proof.
4 Additional comments
1) It follows from a) that any finite set of vertices asymptotically belong to the same tree of the spanning forest.
2) Taking $l=1$, we see that in particular, as $n$ increases to infinity, the distribution of the graph distance of any fixed vertex to the root rescaled by $\sqrt{n}$ converges to the density $xe^{-x^{2}/2}$. Note that the scaling by $\sqrt{n}$ appeared already in [11].
3) If $\kappa=1$, the result can be interpreted as giving the asymptotic distribution of the subtree connecting $l+1$ vertices in the unrooted uniform spanning tree on $K_{n+1}$.
4) Note that for finite $n$, $Q_{L}$ can be non-binary but the corresponding probability tends to 0 as $n\uparrow\infty$.
5) Note that the limiting distribution we get is independent of $\kappa$. On the other hand, it is shown in [8], among other results, that the probability for two vertices to belong to different trees is equivalent to $\frac{\kappa\sqrt{\pi}}{\sqrt{2n}}$ and that the probability for two vertices to be on the same branch starting from the root is equivalent to $\frac{\sqrt{2}\pi}{\sqrt{n}}$. A more general result could be looked for, considering all types of atypical topologies for $Q_{L}$
6) As $l$ increases, we get from b) a consistent family of distributions on $\mathbb{R}_{+}$-extensions of binary tree classes with $l$ leaves. This family is the consistent family of « proper $k$-trees » defined by Aldous in section 4.3 of [3]. It is shown in [3] that this family can be represented by a « Brownian continuum random tree »0 (CRT) as it satisfies a tightness condition. Other constructions of this CRT are given in [2], [1], using Brownian excursions or branching processes. See also [9], [10], [6].