### v5

# Biased spanning forests and Gibbs partitions

Abstract
We study biased spanning forests on the complete graph $K_{N}$, obtained by assigning weight proportional to $\kappa^{r}$ to spanning trees on $K_{N}\cup\{\Delta\}$, where $r$ is the degree of a distinguished root $\Delta$.
Fixing a finite set $L$ of marked vertices, we analyse the minimal $\Delta$-rooted subtree connecting $L$ to $\Delta$. In this sense, we investigate the effect of partially sampling a large random spanning forest through finitely many vertices.
For fixed $\kappa$, the reduced subtree is asymptotically a uniformly distributed binary tree, and graph distances rescaled by $\sqrt{N}$ converge jointly in distribution to an explicit limit introduced by Aldous in his study of the Brownian CRT (continuum random tree).
We show that the scale $\kappa\asymp\sqrt{N}$ is critical: if $\kappa=o(\sqrt{N})$, the marked vertices lie asymptotically in a single component; if $\kappa\gg\sqrt{N}$, the induced partition is asymptotically discrete and the distances to $\Delta$ are negligible on the $\sqrt{N}$ scale.
In the critical regime $\kappa=c\sqrt{N}$, the induced partition of the marked vertices converges to a non-degenerate limit law. This limit is an explicit Gibbs partition of type $\alpha=\tfrac{1}{2}$, whose coefficients admit a closed one-dimensional integral representation. We also note that the limiting EPPF we obtain coincides with a conditional excursion-partition law for reflected Brownian bridges identified by Pitman.
Our approach remains entirely discrete and combinatorial: all asymptotics are obtained from explicit determinant formulas and elementary expansions, without invoking continuum limits.
1 Introduction
Given a random spanning tree $\Upsilon$ of the complete graph $K_{N}$ with $N$ vertices, and a fixed subset $L=\{a_{1},\dots,a_{l}\}$ of $l<N$ labelled vertices, a natural question is the following: how are these $l$ vertices connected within the tree when $N$ is large? More precisely, considering the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ containing $L$, what is the asymptotic distribution, as $N\to\infty$, of its topological type and of the graph distances between its leaves and internal nodes?
This problem can be viewed as a finite-dimensional description of the local geometry of large random spanning trees.
Spanning forests with killing.
We formulate the problem in a slightly more general setting. We distinguish one vertex $\Delta$ as a root and consider random spanning trees on $K_{N}\cup\{\Delta\}$ with weight proportional to $\kappa^{r}$, where $\kappa>0$ is a parameter and $r$ denotes the degree of $\Delta$. Equivalently, when restricted to $K_{N}$, this induces a $\kappa$-biased spanning forest in which a forest with $r$ components has weight proportional to $\kappa^{\,r-1}$.
When $\kappa$ is fixed, this model differs only mildly from the uniform spanning tree. However, when $\kappa$ varies with $N$, qualitatively new phenomena appear. In particular, the scale $\kappa\asymp\sqrt{N}$ turns out to be critical. Our approach remains entirely discrete and combinatorial: all asymptotics are obtained from explicit determinant formulas and elementary expansions, without invoking continuum limits.
Main mechanism.
The asymptotic analysis throughout the paper follows a single principle. For fixed reduced tree shapes and fixed integer edge-length vectors, the Green-determinant formula yields an explicit finite-$N$ expression. After a suitable $\sqrt{N}$ rescaling of the edge lengths, this expression admits a limit. Summing these limits over all integer length vectors and all binary reduced shapes produces Riemann sums which converge to explicit integrals. A crucial point is that the sum of these integrals equal $1$. This shows that asymptotically no probability mass is lost outside the $\sqrt{N}$ scaling window and outside the class of binary reduced trees. Binary shapes therefore emerge with total limiting mass $1$, and their distribution becomes uniform by symmetry.
Fixed-$\kappa$ regime.
In Section 2 we show that, for fixed $\kappa$ and fixed $m$, the value of $\kappa$ is asymptotically irrelevant for the local structure of $\Upsilon_{L}$: the reduced tree of $\Upsilon_{L}$ is asymptotically uniformly distributed among binary trees with leaves $L$, and the graph distances, rescaled by $\sqrt{N}$, converge jointly to an explicit distribution. As $l$ varies, these distributions form a consistent family.
This family coincides with the consistent family of “proper $k$-trees” introduced by Aldous in his study of the Brownian CRT (continuum random tree). In contrast, the present paper derives these asymptotic laws purely by discrete enumeration and elementary asymptotics, without appealing to continuum constructions.
Critical regime and phase transition.
A different phenomenon occurs when
$\kappa=c\sqrt{N},\qquad c>0.$
This scale is critical. If $\kappa=o(\sqrt{N})$, the induced partition of $L$ is asymptotically trivial: with high probability all marked vertices lie in a single component. If $\kappa\gg\sqrt{N}$, the opposite triviality holds: the induced partition is asymptotically the discrete partition, and the graph distance from each marked vertex to $\Delta$ is $o(\sqrt{N})$. Only in the critical regime $\kappa\sim c\sqrt{N}$ do these two effects balance, producing a non-degenerate limiting partition.
In Section 3 we show that in this regime the reduced object $Q_{L}$ obtained from $\Upsilon_{L}$ is, with high probability, a bouquet consisting of $r\in\{1,\dots,l\}$ binary trees attached to $\Delta$. Moreover, for each fixed $r$ and each fixed bouquet configuration (i.e. a partition $\pi$ of $L$ into $r$ blocks, together with a binary tree on each block), all such configurations have the same limiting probability, namely
$\operatorname{\mathbb{P}}\bigl(Q_{L}=\text{given bouquet with $r$ trees}\bigr)\longrightarrow I_{l,r}(c),$
[…]
$\operatorname{\mathbb{P}}\bigl(|\Pi_{N}|=r\bigr)\longrightarrow C_{l,r}\,I_{l,r}(c),$
where $|\Pi_{N}|$ is the number of blocks of the induced partition of $L$ and $C_{l,r}$ is the number of binary forests on $l$ labelled leaves with $r$ components. Conditionally on $|\Pi_{N}|=r$, the internal subtrees defined by the blocks remain asymptotically independent, following the same binary-tree limit law as in the fixed-$\kappa$ case.
Explicit Gibbs structure.
We show in section 3 that the limiting partition in the regime $\kappa\sim c\sqrt{N}$ is Gibbs of type $\alpha=\tfrac{1}{2}$ in the sense of Gnedin and Pitman [8], with block weights $w_{m}=(2m-3)!!$ and explicitly computable coefficients $V_{l,r}(c)$. These weights are normalized by the identity $\sum_{r=1}^{l}C_{l,r}\,I_{l,r}(c)=1$, proved in Section 3. This family is not of Poisson–Dirichlet type, and arises here from a concrete combinatorial model. It appears that the limiting EPPF in the critical regime agrees with Pitman’s conditional excursion partition for a reflected Brownian bridge given its local time at 0.
Dyck paths and excursion partitions.
This latter fact is clearly related to the existence of a path-based encoding of the tree $\Upsilon_{L}$. Once an order is fixed on the leaves (for instance the LL–order used throughout), the $\Delta$–rooted tree $\Upsilon_{L}$ admits a canonical contour/height exploration, whose height process is a Dyck path of length twice the number of edges. Under this encoding, the blocks of the induced partition of the marked vertices correspond to excursion intervals of the height process.
Multitype extension
A multi-cemetery extension can be treated similarly: Wilson’s algorithm can be run in such a way that whenever a loop-erased walk is killed, one records which cemetery point was used. This produces a spanning forest whose components are coloured by cemetery labels, but we focus on the single-root case as ”multitype” versions of the construction and of the asymptotics study require no new ideas.
The relation with previous work is discussed in the final section.
2 Spanning trees in the complete graph
2.1 Definitions, first properties and statement
Green determinants.
Consider a graph $\mathcal{G}$ equipped with conductances and a non-vanishing killing function defined on the set of vertices, here denoted by $X$. These data allow to define a Green matrix $G$ indexed by pairs of vertices. Adding a cemetery point $\Delta$, we can consider the value of the killing function at any vertex $x$ as a conductance between $x$ and $\Delta$. Then a spanning forest of rooted trees on $\mathcal{G}$ can be identified to A spanning tree on the extended graph $\mathcal{G}_{\Delta}$, rooted in $\Delta$ induces a spanning forest on $\mathcal{G}$. An extension of Cayley’s theorem (cf. for example [11], section 8-2 in [12]) shows that if we define the weight of such a spanning tree to be the product of the conductances of its edges, the sum of these weights is the inverse of the determinant of $G$. It provides naturally a probability $\mathbb{P}$ on spanning trees rooted in $\Delta$ (and consequently a probability on spanning forests.)
This applies in particular to the complete graph $K_{N}$ with vertices $\{1,2,...,N\}$ endowed with unit conductances and a constant killing function $\kappa$. It is naturally associated with a random walk for which at each step, the random walk is killed with probability $\frac{\kappa}{\kappa+N-1}$.
The corresponding probability on $\Delta$-rooted spanning trees will be denoted by $\mathbb{P}^{(N),\kappa}$. If $\kappa=1$, the spanning tree rooted in $\Delta$ can be identified with a uniform spanning tree on $K_{N+1}$. Local limits for these objects have been determined in [6], extending a result of Grimmett [9].
The expression of the Green matrix $G^{(N),\kappa}$ is:
$\frac{1}{\kappa(N+\kappa)}(\kappa I+J)$
where $J$ denotes the $(N,N)$ matrix with all entries equal to $1$. It is easy to check that for any $(d,d)$ square matrix $M$ with diagonal entries equal to $a+b$ and off diagonal entries equal to $b$, $\det(M)=a^{d-1}(a+db)$. Hence we get the following identity:
$\det(G^{(N),\kappa})=\frac{1}{\kappa(N+\kappa)^{N-1}}$ (1)
Moreover, for $0<d<N$,
$\det(G^{(N),\kappa}_{i,j},\,1\leq i,j\leq d)=\frac{d+\kappa}{\kappa(N+\kappa)^{d}}$ (2)
Wilson’s algorithm.
Coming back to the general case, a simple way to sample a spanning tree $\Upsilon$ under the probability $\mathbb{P}$ is to perform Wilson’s algorithm (cf. [23]) which is based on loop erasure. The probability of a loop-erased path $\xi$ from a given vertex $\xi_{0}$ to $\Delta$ is the product of its conductances multiplied by the determinant of the restriction of the Green matrix to the vertices of $\xi$. Wilson’s algorithm starts by choosing any order on the vertices and constructing a loop erased path from the first vertex to $\Delta$, then at each step it constructs a loop erased path from the first unused vertex to the tree of used vertices (including $\Delta$) until all vertices have been used.
We are interested in the smallest subtree $\Upsilon_{L}$ of $\Upsilon$ connecting a set $L$ of $l$ vertices and $\Delta$. For that, we can run Wilson algorithm after choosing an order in which these vertices are the first $l$ vertices. Given a tree $Y$ imbedded in the graph and rooted in $\Delta$ containing $L$ and whose set of leaves is included in $L$, the spanning tree $\Upsilon$ produced by the algorithm will contain it iff after the $l$-th step of the algorithm the tree of used vertices is exactly $Y$. Then we will have $\Upsilon_{L}=Y$.
It is easy to check from the proof of Wilson’s algorithm given in [12] that for any tree $Y$ rooted in $\Delta$, the probability $\mathbb{P}(Y=\Upsilon_{L})$ is given by the product of the conductances of the edges of $Y$ multiplied by the determinant of the restriction of the Green matrix to the vertices of $Y\setminus\{\Delta\}$.
Consequently, in the case of the complete graph with $N$ vertices $\{1,...,N\}$ endowed with unit conductances and a constant killing factor $\kappa$, if $Y$ has $d$ vertices (root excepted) and $r$ edges incident to $\Delta$, by formula 2, we have:
Lemma 2.1.
$\mathbb{P}^{(N),\kappa}(Y=\Upsilon_{L})=\frac{\kappa^{r-1}(d+\kappa)}{(N+\kappa)^{d}}$
Reduced trees and lengths.
A spanning tree defines a map $p$ from $X$ into $X\cup\{\Delta\}$ fixing $\Delta$ and such that for any vertex $x$ in $X$ and positive integer $m$, $p^{m}(x)\neq x$. If $p(y)=x$, we say that $y$ is a child of $x$. If $p^{m}(y)=x$ for some non-negative (positive) integer $m$, we say that $y$ is above (strictly above) $x$ and that $x$ is below (strictly below) $y$. A node is by definition a vertex with at least two children. These notions are extended to any tree imbedded in the graph and rooted in $\Delta$ when we replace $X$ by the set of tree vertices.
Let us come back to the case of the complete graph $K_{N}$.
We say that a tree imbedded in the complete graph and rooted in $\Delta$ is a $L$-tree if it contains $L$ and his set of leaves, denoted by $L^{*}$, is contained in $L$. $\Upsilon_{L}$ is always a $L$-tree. We denote by $\mathcal{N}$ the set of nodes of $\Upsilon_{L}$.
There is a mapping $j$ from $\mathcal{N}\cup L$ into parts of $L$. It maps a vertex $x$ to the set of elements of $L$ which are above it in the tree. In particular, it sends a leaf $i$ into $\{i\}$. We also set $j(\Delta)=L$.
We say that a tree rooted in $\Delta$ is $L$-reduced if all its vertices which are not elements of $L$ or nodes have been removed. A $L$-tree can be decomposed into its $L$-reduction which is still a $L$ -tree and finite sequences of intermediate inner vertices with a single child.
It is clearly enough to consider the case $L=\{1,...l\}$. Then the natural order on $L$ is used to define an order on the children of each internal node: we order these children by the minimum element of the set of leaves below them. (In the binary case this simply fixes a canonical left/right orientation at each branch point.) This turns $Y$ into a plane rooted tree in a deterministic way.
Perform the depth-first contour traversal of the resulting plane rooted tree, starting and ending at $\Delta$; this visits each edge exactly twice and defines an order on all vertices of $Y$.
In order to capture the way elements of $L$ are connected, in other words the geometry of the subtree connecting them, only the labels of $L$ do matter. This leads us to define an equivalence relation: we say that two $L$-trees are equivalent if they can be exchanged by a permutation of $\{l+1,...,N\}$. Note that equivalence preserves the $L$-order we defined on $\mathcal{N}\cup L$. The (equivalence) class of the $L$-reduction of $\Upsilon_{L}$ denoted by $Q_{L}$ inherits a tree structure. Its inner vertices are the images of $\mathcal{N}\cup L\setminus L^{*}$ by $j$ and its leaves singletons in $L^{*}$. The class of $\Upsilon_{L}$ is determined by $Q_{L}$ and the numbers of intermediate inner vertices between each vertex in $\mathcal{N}\cup L$ and the first vertex in $\mathcal{N}\cup L\setminus L^{*}\cup\{\Delta\}$ below it, hence by a map $u_{L}$ from the set of vertices of $Q_{L}$ into the set $\mathbb{N}$ of nonnegative integers. We can say that $(Q_{L},u_{L})$ is a $\mathbb{N}$-extension of $Q_{L}$.
We say that a $L$-tree is binary iff $L^{*}=L$, every internal vertex has exactly two children and the root only one. An easy induction shows that the set $\mathcal{B}_{l}$ of binary $L$-tree classes has cardinality $c_{l}=\prod_{i=1}^{l-1}(2i-1)=\frac{(2l-3)!}{2^{l-2}(l-2)!}$ and that these trees have l1l-1 internal vertices. Indeed, adding one leaf is done by choosing a vertex, adding a node just below it and connecting it to the new leaf. Consider now binary tree extensions: using the $L$-order on leaves and nodes, $u_{L}$ becomes a (2l1)(2l-1)-tuple of nonnegative integers $(u_{L}(i),1\leq i\leq 2l-1)$.
Contour (Dyck-path) encoding.
We recall a standard encoding of rooted trees by walks on $\operatorname{\mathbb{Z}}_{\geq 0}$. A Dyck path of length $2n$ is a function $H:\{0,1,\dots,2n\}\to\operatorname{\mathbb{Z}}_{\geq 0}$ such that $H(0)=H(2n)=0$ and $H(k+1)-H(k)\in\{+1,-1\}$ for all $k$.
Let $v_{k}$ be the visited vertex at step $k$ by the contour path. If $Y$ is a rooted tree with $n$ vertices, denoting by $d_{Y}$ the graph distance in $Y$, set
$H_{Y}(k):=d_{Y}(\Delta,v_{k}),\qquad k=0,1,\dots,2n,$
Then $H_{Y}$ is a Dyck path of length $2n$.
[…]
Asymptotics.
We can now formulate precisely our problem which is to show that as $N$ increases to infinity, the probability that the reduction of $\Upsilon_{L}$ is binary converges to one, and to determine the asymptotic behavior of the joint distribution of the pair $(Q_{L},u_{L})$. This is given in the following:
Theorem 2.1.
a) $\mathbb{P}^{(N),\kappa}(L^{*}=L)$ converges to 1 as $N\uparrow\infty$.
b) Given $\alpha\in\mathcal{B}_{l}$,
$\lim_{N\uparrow\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)=c_{l}^{-1}.$
i.e. the distribution of $Q_{L}$ converges towards the uniform distribution on $\mathcal{B}_{l}$.
c) Given $t\in\mathbb{R}_{+}^{2l-1}$,
$\lim_{N\uparrow\infty}N^{(2l-1)/2}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{N}])=(\sum_{1}^{2l-1}t_{i})e^{-(\sum_{1}^{2l-1}t_{i})^{2}/2}$
in which $t\sqrt{N}]$ denotes the 2l12l-1-tuple of integers $([\sqrt{N}t_{1}],...,[\sqrt{N}t_{2l-1}])$.
2.2 Proof
Note that a) follows directly from b) since there are $c_{l}$ binary trees with $l$ labeled leaves.
Then we start with the proof of c). Note that setting $\Sigma=\sum_{1}^{2l-1}[t_{i}\sqrt{N}]$ there are $\prod_{i=0}^{\Sigma-1}(N-l-i)$ possible choices for the inner vertices of $\Upsilon_{L}$, given that $u_{L}=[t\sqrt{N}]$. From lemma 2.1, we get that
$\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=[t\sqrt{N}])=\frac{\Sigma+l+\kappa}{(N+\kappa)^{\Sigma+l}}\prod_{i=0}^{\Sigma-1}(N-l-i).$
(with the usual convention that an empty product equals 1).
We define
$L_{N}:=\log(\frac{\Sigma+l+\kappa}{(N+\kappa)^{\Sigma+l}}\prod_{i=0}^{\Sigma-1}(N-l-i)),$
set $\Sigma=\sigma\sqrt{N}$, and assume that $0\leq\sigma\leq\sigma_{\max}$, with $l$ and $\kappa$ fixed.
We prove that there exist constants $C>0$ and $N_{0}$ (depending only on $\sigma_{\max},l,\kappa$) such that for all $N\geq N_{0}$ and all $\sigma\in[0,\sigma_{\max}]$,
$\left|\,L_{N}-\Bigl(\log(\Sigma+l+\kappa)-l\log N-\frac{\sigma^{2}}{2}\Bigr)\right|\leq\frac{C}{\sqrt{N}}.$
Uniform estimate for the critical scaling
Recall
$L_{N}=\log(\Sigma+l+\kappa)-(\Sigma+l)\log(N+\kappa)+\sum_{i=0}^{\Sigma-1}\log(N-l-i),$
with $\Sigma=\sum_{i=1}^{2l-1}k_{i}$ and $\Sigma\leq\sigma_{\max}\sqrt{N}$.
Rewrite
$\log(N+\kappa)=\log N+\log\!\left(1+\frac{\kappa}{N}\right),\qquad\log(N-l-i)=\log N+\log\!\left(1-\frac{l+i}{N}\right),$
so that
$L_{N}=\log(\Sigma+l+\kappa)-l\log N-(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)+\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right).$
Choose $N_{0}$ such that for all $N\geq N_{0}$ and $\Sigma\leq\sigma_{\max}\sqrt{N}$,
$0\leq\frac{\kappa}{N}\leq\frac{1}{2},\qquad 0\leq\frac{l+\Sigma-1}{N}\leq\frac{1}{2}.$
(1) The term $\log(1+\kappa/N)$. For $0\leq y\leq 1/2$,
$0\leq\log(1+y)\leq y+y^{2}.$
Hence
$\left|(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)\right|\leq(\Sigma+l)\left(\frac{\kappa}{N}+\frac{\kappa^{2}}{N^{2}}\right).$
Using $\Sigma\leq\sigma_{\max}\sqrt{N}$,
$\left|(\Sigma+l)\log\!\left(1+\frac{\kappa}{N}\right)\right|\leq\frac{\kappa(\sigma_{\max}+l)}{\sqrt{N}}+\frac{\kappa^{2}(\sigma_{\max}+l)}{N^{3/2}}\leq\frac{C_{1}}{\sqrt{N}},$
where $C_{1}=\kappa(\sigma_{\max}+l)+\kappa^{2}(\sigma_{\max}+l)$.
(2) The product term. For $0\leq x\leq 1/2$,
$\log(1-x)=-x+\varepsilon(x),\qquad|\varepsilon(x)|\leq\frac{x^{2}}{1-x}\leq 2x^{2}.$
Set $x_{i}=(l+i)/N$. Then
$\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right)=-\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)+\sum_{i=0}^{\Sigma-1}\varepsilon(x_{i}).$
Moreover,
$\left|\sum_{i=0}^{\Sigma-1}\varepsilon(x_{i})\right|\leq 2\sum_{i=0}^{\Sigma-1}\left(\frac{l+i}{N}\right)^{2}\leq\frac{2}{N^{2}}\,\Sigma\,(l+\Sigma)^{2}\leq\frac{C_{2}}{\sqrt{N}},$
with $C_{2}=2\sigma_{\max}(l+\sigma_{\max})^{2}$.
Now
$\sum_{i=0}^{\Sigma-1}(l+i)=\Sigma l+\frac{\Sigma(\Sigma-1)}{2},$
so
$\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)=\frac{\Sigma^{2}}{2N}+\frac{\Sigma l}{N}-\frac{\Sigma}{2N}.$
Hence
$\left|\frac{1}{N}\sum_{i=0}^{\Sigma-1}(l+i)-\left(\frac{\Sigma^{2}}{2N}+\frac{\Sigma l}{N}\right)\right|=\frac{\Sigma}{2N}\leq\frac{\sigma_{\max}}{2\sqrt{N}}=:\frac{C_{3}}{\sqrt{N}}.$
Combining,
$\sum_{i=0}^{\Sigma-1}\log\!\left(1-\frac{l+i}{N}\right)=-\frac{\Sigma^{2}}{2N}-\frac{\Sigma l}{N}+r_{N},\qquad|r_{N}|\leq\frac{C_{2}+C_{3}}{\sqrt{N}}.$
(3) Conclusion. Inserting (1) and (2) into $L_{N}$ yields
$\left|L_{N}-\left(\log(\Sigma+l+\kappa)-l\log N-\frac{\Sigma^{2}}{2N}\right)\right|\leq\frac{C}{\sqrt{N}},$
with $C=C_{1}+C_{2}+C_{3}+l\,\sigma_{\max}$, uniformly for $\Sigma\leq\sigma_{\max}\sqrt{N}$.
Write
$\delta_{N}:=L_{N}-\Bigl(\log(\Sigma+l+\kappa)-l\log N-\frac{\Sigma^{2}}{2N}\Bigr),\qquad|\delta_{N}|\leq\frac{C}{\sqrt{N}}.$
Then
$e^{L_{N}}=(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)}\,e^{\delta_{N}}.$
For $N$ large enough so that $C/\sqrt{N}\leq 1$, we have $|e^{\delta_{N}}-1|\leq 2|\delta_{N}|$, hence
$\left|e^{L_{N}}-(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)}\right|\leq\frac{2C}{\sqrt{N}}\,(\Sigma+l+\kappa)\,N^{-l}e^{-\Sigma^{2}/(2N)},$
uniformly for all $\sigma\in[0,\sigma_{\max}]$. This yields the uniform estimate in c). Moreover, for all $N$ sufficiently large and all $k\in\mathbb{N}^{2l-1}$ satisfying $k_{i}\leq\sigma_{\max}\sqrt{N}$,
$\left|N^{(2l-1)/2}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=k)-\left(\sum_{i=1}^{2l-1}\frac{k_{i}}{\sqrt{N}}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}k_{i})^{2}}{2N}\right)\right|\leq\frac{C}{\sqrt{N}}.$
To prove b), we first compute the normalising integral:
$\int_{\mathbb{R}_{+}^{2l-1}}\left(\sum_{i=1}^{2l-1}t_{i}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}t_{i})^{2}}{2}\right)dt=c_{l}^{-1}.$
Fix $\sigma_{\max}>0$. By the uniform estimate obtained in part (c),
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)$ $\geq\liminf_{N\to\infty}\sum_{\begin{subarray}{c}k\in\mathbb{N}^{2l-1}\\ k_{i}\leq\sigma_{\max}\sqrt{N}\end{subarray}}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha,u_{L}=k)$
$=\lim_{N\to\infty}N^{-(2l-1)/2}\sum_{\begin{subarray}{c}k\in\mathbb{N}^{2l-1}\\ k_{i}\leq\sigma_{\max}\sqrt{N}\end{subarray}}\left(\sum_{i=1}^{2l-1}\frac{k_{i}}{\sqrt{N}}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}k_{i})^{2}}{2N}\right).$
The last sum is precisely the Riemann sum over the cube $0,\sigma_{\max}]^{2l-1}$ with mesh size $h_{N}=1/\sqrt{N}$ of the continuous function
$f(t)=\left(\sum_{i=1}^{2l-1}t_{i}\right)\exp\!\left(-\frac{(\sum_{i=1}^{2l-1}t_{i})^{2}}{2}\right).$
Therefore,
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)\geq\int_{[0,\sigma_{\max}]^{2l-1}}f(t)\,dt.$
Letting $\sigma_{\max}\to\infty$, monotone convergence gives
$\liminf_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)\geq\int_{\mathbb{R}_{+}^{2l-1}}f(t)\,dt=c_{l}^{-1}.$
Since the probabilities $\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)$ sum to $1$ over the $c_{l}$ binary shapes, we must have equality:
$\lim_{N\to\infty}\mathbb{P}^{(N),\kappa}(Q_{L}=\alpha)=c_{l}^{-1}.$
This completes the proof of b).
2.3 Additional comments
1) It follows from a) that any finite set of vertices asymptotically belong to the same tree of the spanning forest.
2) Taking $l=1$, we see that in particular, as $N$ increases to infinity, the distribution of the graph distance of any fixed vertex to the root rescaled by $\sqrt{N}$ converges to the density $xe^{-x^{2}/2}$. Note that the scaling by $\sqrt{N}$ appeared already in [22].
3) If $\kappa=1$, the result can be interpreted as giving the asymptotic distribution of the subtree connecting $l+1$ vertices in the unrooted uniform spanning tree on $K_{N+1}$.
4) Note that for finite $n$, $Q_{L}$ can be non-binary but the corresponding probability tends to 0 as $\uparrow\infty$.
5) Note that the limiting distribution we get is independent of $\kappa$. On the other hand, it is shown in [12], among other results, that the probability for two vertices to belong to different trees is equivalent to $\frac{\kappa\sqrt{\pi}}{\sqrt{2n}}$ and that the probability for two vertices to be on the same branch starting from the root is equivalent to $\frac{\sqrt{2}\pi}{\sqrt{n}}$. A more general result could be looked for, considering all types of atypical topologies for $Q_{L}$
3 The critical case: $\kappa=c\sqrt{N}$
We now consider the case where the killing parameter depends on $N$ and satisfies
$\kappa=\kappa_{N}=c\sqrt{N},\qquad c>0\ \text{fixed}.$
3.1 Statement of the main result
In contrast with the previous regime (where $\kappa$ is fixed), we will prove that as $N\uparrow\infty$ the subtree $\Upsilon_{L}$ connecting $L$ to $\Delta$ may now be a bouquet of several disjoint binary trees rooted in $\Delta$. We say that a $L$-tree is a bouquet of $r$ binary trees iff $L^{*}=L$, every internal vertex has exactly two children and the root exactly $r$.
Let $\pi=\{B_{1},\dots,B_{r}\}$ be a partition of $L=\{1,\dots,l\}$ into $r$ non–empty blocks. For each block $B_{j}$ (with $|B_{j}|=l_{j}$), let $\alpha_{j}\in\mathcal{B}_{l_{j}}$ be a binary $B_{j}$–tree class. The configurations we consider are the bouquets of $r$ binary $B_{j}$-trees rooted in $\Delta$. Their leaves are exactly $L$. The corresponding $Q_{L}$’s can then be identified with a $r$-tuple $(\alpha_{j},\;j=1,...r)$ of $L$- reduced binary trees classes.
As in the case $r=1$, to fully determine the class of $\Upsilon_{L}$, the $Q_{L}$’s are completed by $\mathbb{N}$–extensions: if $u=(u_{i})_{1\leq i\leq d}$ denotes the numbers of intermediate vertices along the edges of the reduced tree, we have:
$\sum_{j=1}^{r}(2l_{j}-1)=2l-r,\qquad\Sigma=\sum_{i=1}^{2l-r}u_{i},\qquad a=l+\Sigma.$
By Lemma 2.1, the probability of a given configuration $(\pi,(\alpha_{j})_{1\leq j\leq r},u)$ is
$\mathbb{P}^{(N),\kappa}(\pi,(\alpha_{j}),u)=\frac{\kappa^{r-1}(l+\Sigma+\kappa)}{(N+\kappa)^{l+\Sigma}}\prod_{i=0}^{\Sigma-1}(N-l-i).$ (3)
An easy recursion shows that
$C_{l,r}=\sum_{\pi\in\mathcal{P}_{l,r}}\prod_{B\in\pi}c_{|B|}$
is the number of classes of bouquets of $r$ binary trees whose leaves are $L$ ($\mathcal{P}_{l,r}$ denotes the set of partitions of $L$ into $r$ blocks).
Theorem 3.1 (Critical regime $\kappa=c\sqrt{N}$).
Fix $l\geq 1$ and $c>0$, and set $\kappa=\kappa_{N}=c\sqrt{N}$. Let $L=\{1,\dots,l\}$.
(a) Binary reduction. The probability that $\Upsilon_{L}$ is not a binary bouquet rooted at $\Delta$ tends to 0:
(b) Limit law of reduced configurations. Let $\pi=\{B_{1},\dots,B_{r}\}$ be a partition of $L$ into $r$ blocks and let $\alpha_{j}\in\mathcal{B}_{|B_{j}|}$ be reduced binary tree classes on each block, as above. Then
$\operatorname{\mathbb{P}}^{(N),c\sqrt{N}}\!\bigl(Q_{L}=(\pi,(\alpha_{j})_{1\leq j\leq r})\bigr)\longrightarrow I_{l,r}(c),$
where $I_{1,1}=1$ and for $l\geq 2$ $I_{l,r}(c)$ is given by:
$I_{l,r}(c)=\frac{c^{\,r-1}}{(2l-r-1)!}\int_{0}^{\infty}(\sigma+c)\sigma^{2l-r-1}e^{-\frac{\sigma^{2}}{2}-c\sigma}\,d\sigma$ (4)
Or alternatively by
$I_{l,r}(c)=\frac{c^{\,r-1}}{(2l-r-2)!}\int_{0}^{\infty}s^{2l-r-2}\exp\!\Bigl(-\frac{s^{2}}{2}-cs\Bigr)\,ds.$ (5)
In particular, writing
$C_{l,r}=\sum_{\pi\in\mathcal{P}_{l,r}}\ \prod_{B\in\pi}c_{|B|}\qquad(1\leq r\leq l),$
one has
$\operatorname{\mathbb{P}}^{(N),c\sqrt{N}}(|\Pi_{N}|=r)\longrightarrow C_{l,r}\,I_{l,r}(c),$
and conditionally on $|\Pi_{N}|=r$, the reduced configuration $(\pi,(\alpha_{j}))$ is asymptotically uniform over the $C_{l,r}$ possibilities.
[…]
$\lim_{N\to\infty}N^{(2l-r)/2}\,\operatorname{\mathbb{P}}^{(N),c\sqrt{N}}\!\bigl(Q_{L}=(\pi,(\alpha_{j})),\ u=[t\sqrt{N}]\bigr)=c^{\,r-1}(\sigma+c)\,\exp\!\Bigl(-\frac{\sigma^{2}}{2}-c\sigma\Bigr).$
3.2 Proof
Note first that c) implies b) by summing over $u$, and that b) implies a) by summing over all bouquet configurations. We therefore focus on proving c).
Proof of c).
Let $t\in\mathbb{R}_{+}^{d}$ and set $u_{i}=[t_{i}\sqrt{N}]$. Denote $\sigma=\sum_{i=1}^{d}t_{i}$ so that $\Sigma=\sigma\sqrt{N}+\theta_{N}$,   $|\theta_{N}|\leq 1$. Proceeding exactly as in the proof of part c) of Theorem 1 (Taylor expansion of the logarithm and uniform error control on bounded cubes), one obtains:
Lemma. For every $\sigma_{\max}>0$ there exist $C>0$ and $N_{0}$ such that, uniformly for $t\in[0,\sigma_{\max}]^{d}$ and $N\geq N_{0}$,
$\Bigg|N^{d/2}\mathbb{P}^{(N),c\sqrt{N}}\bigl(\pi,(\alpha_{j}),u=[t\sqrt{N}]\bigr)-c^{\,r-1}(\sigma+c)e^{-\frac{\sigma^{2}}{2}-c\sigma}\Bigg|\leq\frac{C}{\sqrt{N}}.$
In particular, for fixed $t$, as $n\uparrow\infty$:
$N^{d/2}\mathbb{P}^{(N),c\sqrt{N}}\bigl(\pi,(\alpha_{j}),u=[t\sqrt{N}]\bigr)\longrightarrow c^{\,r-1}(\sigma+c)e^{-\frac{\sigma^{2}}{2}-c\sigma}.$ (6)
Riemann sums and lower bound.
Fix $\sigma_{\max}>0$. Summing over $u_{i}\leq\sigma_{\max}\sqrt{N}$ and using the uniform estimate obtained earlier, we obtain, as in the proof of Theorem 1,
$\lim_{N\to\infty}\sum_{\begin{subarray}{c}u\in\mathbb{N}^{d}\\ u_{i}\leq\sigma_{\max}\sqrt{N}\end{subarray}}\mathbb{P}^{(N),c\sqrt{N}}(\pi,(\alpha_{j}),u)=\int_{[0,\sigma_{\max}]^{d}}c^{\,r-1}(\sigma+c)e^{-\frac{\sigma^{2}}{2}-c\sigma}\,dt.$
Letting $\sigma_{\max}\to\infty$ and using monotone convergence gives
$\liminf_{N\to\infty}\mathbb{P}^{(N),c\sqrt{N}}(\pi,(\alpha_{j}))\ \geq\ \int_{\mathbb{R}_{+}^{d}}c^{\,r-1}(\sigma+c)e^{-\frac{\sigma^{2}}{2}-c\sigma}\,dt.$ (7)
Since the integrand depends on $t$ only through $\sigma=\sum_{i=1}^{d}t_{i}$, the simplex–volume identity gives
$\int_{\mathbb{R}_{+}^{2l-r}}c^{\,r-1}(\sigma+c)e^{-\frac{\sigma^{2}}{2}-c\sigma}\,dt=\frac{c^{\,r-1}}{(2l-r-1)!}\int_{0}^{\infty}(\sigma+c)\sigma^{2l-r-1}e^{-\frac{\sigma^{2}}{2}-c\sigma}\,d\sigma=I_{l,r}(c).$ (8)
Summing (7) over all partitions $\pi$ with $r$ blocks and all binary choices $(\alpha_{j})$, we obtain that the total limiting mass of configurations with $r$ components is larger or equal to:
$M_{l,r}(c)=C_{l,r}\,I_{l,r}(c).$
Summing over $r=1,\dots,l$, we obtain
$\liminf_{N\to\infty}\sum_{r=1}^{l}\mathbb{P}^{(N),c\sqrt{N}}(\text{$r$ components})\ \geq\ \sum_{r=1}^{l}C_{l,r}\,I_{l,r}(c).$
We now have to prove that
$\sum_{r=1}^{l}C_{l,r}\,I_{l,r}(c).=1.$ (9)
Since probabilities cannot sum to more than $1$, it will follow that all the above inequalities are in fact equalities. In particular that:
$\lim_{N\to\infty}\mathbb{P}^{(N),c\sqrt{N}}(\pi,(\alpha_{j}))=I_{l,r}(c)$
Recursion.
For $n\in\operatorname{\mathbb{N}}$ and $c>0$ define
$A_{n}(c):=\int_{0}^{\infty}\sigma^{n}e^{-\sigma^{2}/2-c\sigma}\,d\sigma.$
Lemma 3.1.
For every $n\geq 1$,
$A_{n+1}(c)=nA_{n-1}(c)-cA_{n}(c).$ (10)
Proof.
Perform an integration by parts over $(0,\infty)$ and check that the boundary term vanishes. ∎
We may rewrite
$I_{l,r}(c)=\frac{c^{r-1}}{(m+1)!}\Big(A_{m+2}(c)+cA_{m+1}(c)\Big),\text{ with }m:=2l-r-2.$ (11)
For $l>1$, applying (10) with $n=m+1$ yields the simplified exact form
$I_{l,r}(c)=\frac{c^{r-1}}{m!}\,A_{m}(c)=\frac{c^{r-1}}{(2l-r-2)!}\int_{0}^{\infty}\sigma^{2l-r-2}e^{-\sigma^{2}/2-c\sigma}\,d\sigma,\qquad(l\geq 2).$ (12)
In the base case $l=1$, $r=1$, and
$I_{1,1}(c)=\int_{0}^{\infty}(\sigma+c)e^{-\sigma^{2}/2-c\sigma}\,d\sigma=\Big[-e^{-\sigma^{2}/2-c\sigma}\Big]_{0}^{\infty}=1.$
Lemma 3.2 (Recursion for $C_{l,r}$).
For $l\geq 2$ and $1\leq r\leq l$,
$C_{l,r}=(2l-r-2)\,C_{l-1,r}+C_{l-1,r-1},$ (13)
with boundary conditions $C_{1,1}=1$ and $C_{l,r}=0$ if $r<1$ or $r>l$.
Proof.
We construct a weighted partition of $L=\{1,\dots,l\}$ into $r$ blocks by inserting the new leaf $l$ into a weighted partition of $L^{\prime}=\{1,\dots,l-1\}$.
Case 1: $l$ forms a singleton block. Start from a partition of $L^{\prime}$ into r1r-1 blocks, with weights $\prod_{B}c_{|B|}$, and add the singleton $\{l\}$. Since $c_{1}=1$, this contributes $C_{l-1,r-1}$.
Case 2: $l$ is inserted into an existing block. Start from a partition $\pi^{\prime}$ of $L^{\prime}$ into $r$ blocks. Choose a block $B$ of size $m$ and enlarge it to size $m+1$. At the level of binary shapes,
$c_{m+1}=(2m-1)c_{m},$
so enlarging the chosen block multiplies its weight by (2m1)(2m-1). Summing (2|B|1)(2|B|-1) over the blocks of $\pi^{\prime}$ gives
$\sum_{B\in\pi^{\prime}}(2|B|-1)=2(l-1)-r=2l-r-2,$
hence the total contribution of Case 2 is $(2l-r-2)\,C_{l-1,r}$. Adding the two cases yields (13). ∎
Define
$S_{l}(c):=\sum_{r=1}^{l}C_{l,r}\,I_{l,r}(c).$
We finally prove that $S_{l}(c)=1$ for all $l\geq 1$.
Since $C_{1,1}=1$, we obtain $S_{1}(c)=1$.
Induction step. Fix $l\geq 2$. Using (13) and reindexing, we obtain
$S_{l}(c)=\sum_{r=1}^{l-1}C_{l-1,r}\Big((2l-r-2)\,I_{l,r}(c)+I_{l,r+1}(c)\Big).$ (14)
We first prove the identity
$(2l-r-2)\,I_{l,r}(c)+I_{l,r+1}(c)=I_{l-1,r}(c),\qquad 1\leq r\leq l-1.$ (15)
Let $m:=2l-r-2\geq 1$. Using (12) for $(l,r)$ and $(l,r+1)$,
$(2l-r-2)I_{l,r}(c)+I_{l,r+1}(c)=m\frac{c^{r-1}}{m!}A_{m}(c)+\frac{c^{r}}{(m-1)!}A_{m-1}(c)=\frac{c^{r-1}}{(m-1)!}\Big(A_{m}(c)+cA_{m-1}(c)\Big).$
For $m\geq 2$, applying (10) with n=m1n=m-1 gives $A_{m}(c)+cA_{m-1}(c)=(m-1)A_{m-2}(c)$, hence
$(2l-r-2)I_{l,r}(c)+I_{l,r+1}(c)=\frac{c^{r-1}}{(m-2)!}A_{m-2}(c).$
Since m2=2(l1)r2m-2=2(l-1)-r-2, this equals $I_{l-1,r}(c)$ by (12), proving (15) for $m\geq 2$. The case $m=1$ can be checked directly.
Plugging (15) into (14) yields
$S_{l}(c)=\sum_{r=1}^{l-1}C_{l-1,r}\,I_{l-1,r}(c)=S_{l-1}(c).$
Therefore $S_{l}(c)=S_{1}(c)=1$ for all $l\geq 1$, which is (9). ∎
3.3 Gibbs structure and a tree-restaurant construction
We briefly recall the notion of Gibbs partitions in the sense of Gnedin and Pitman [8].
Let $\Pi_{n}$ be a random partition of $\{1,\dots,n\}$. It is called exchangeable if its law is invariant under permutations of the labels. Such a partition is characterised by its exchangeable partition probability function (EPPF)
$p(n_{1},\dots,n_{r})=\operatorname{\mathbb{P}}(\Pi_{n}\text{ has }r\text{ blocks of sizes }n_{1},\dots,n_{r}),$
where $n_{1},\dots,n_{r}\geq 1$ and $n_{1}+\cdots+n_{r}=n$. By exchangeability, $p$ depends only on the block sizes, not on the order of the blocks.
Gibbs partitions.
Fix a parameter $\alpha<1$. An exchangeable partition is said to be Gibbs of type $\alpha$ if its EPPF factorises as
$p(n_{1},\dots,n_{r})=V_{n,r}\prod_{i=1}^{r}w_{n_{i}},$ (16)
where
$w_{m}=(1-\alpha)_{m-1}=(1-\alpha)(2-\alpha)\cdots(m-1-\alpha),\qquad m\geq 1,$
and $(x)_{k}$ denotes the rising factorial. The coefficients $V_{n,r}$ are nonnegative numbers satisfying the recursion
$V_{n,r}=(n-\alpha r)\,V_{n+1,r}+V_{n+1,r+1},$ (17)
together with the normalization
$\sum_{r=1}^{n}V_{n,r}\!\!\sum_{\begin{subarray}{c}n_{1}+\cdots+n_{r}=n\\ n_{i}\geq 1\end{subarray}}\prod_{i=1}^{r}w_{n_{i}}=1.$
The best-known Gibbs partitions are the two-parameter Poisson–Dirichlet distributions $PD(\alpha,\theta)$, for which
$V_{n,r}^{PD}=\frac{(\theta+\alpha)_{r-1;\alpha}}{(\theta+1)_{n-1}},$
where
$(x)_{k;\alpha}=x(x+\alpha)\cdots(x+(k-1)\alpha)$
is the generalized rising factorial. In this case the probability of creating a new block depends only on $(n,r)$ through a linear expression.
In our setting, the limiting partition in the critical regime $\kappa=c\sqrt{N}$ is Gibbs of type $\alpha=\tfrac{1}{2}$. The block weights are
$w_{m}=(1-\tfrac{1}{2})_{m-1}=\frac{(2m-3)!!}{2^{\,m-1}},$
The coefficients $V_{l,r}(c)$ are given explicitly by
$V_{l,r}(c)=2^{\,l-r}\,I_{l,r}(c),$
where $I_{l,r}(c)$ admits the one-dimensional integral representation derived in Section 3. Unlike the Poisson–Dirichlet case, these coefficients do not reduce to a simple ratio of factorials and depend non-linearly on $c$.
Sequential construction
This Gibbs($1/2$) structure admits a natural sequential description. Suppose that after inserting $l$ leaves the partition of $\{1,\dots,l\}$ has $r$ blocks of sizes $n_{1},\dots,n_{r}$. To insert leaf $l+1$:
  • (i)
    Choose an existing block $i$ with probability
    $\frac{(n_{i}-\tfrac{1}{2})V_{l+1,r}(c)}{V_{l,r}(c)}.$
    and attach the new leaf uniformly to one of the $2n_{i}-1$ possible insertion sites in the corresponding binary tree.
  • (ii)
    Create a new singleton block with probability
    $\theta_{l,r}(c)=\frac{V_{l+1,r+1}(c)}{V_{l,r}(c)}.$
Since $\sum_{i}(n_{i}-\tfrac{1}{2})=l-\tfrac{r}{2}$, this defines a valid transition rule. As induction shows that this construction reproduces exactly the weights $V_{l,r}(c)\prod_{B\in\pi}w_{|B|}$.
Remark 3.1.
The explicit formula shows that
$\theta_{l,r}(c)\to 0\quad(c\downarrow 0),\qquad\theta_{l,r}(c)\to 1\quad(c\uparrow\infty),$
in agreement with the subcritical and supercritical regimes described below in section 3.4.
Remark 3.2.
A mixture identity Fix $\beta>-1$ and let $C$ be a positive random variable with density
$h_{\beta}(c)=\frac{1}{Z_{\beta}}\,c^{\beta}e^{-c^{2}/2}\,\mathbf{1}_{\{c>0\}},\qquad Z_{\beta}=\int_{0}^{\infty}c^{\beta}e^{-c^{2}/2}\,dc=2^{(\beta-1)/2}\Gamma\!\Big(\frac{\beta+1}{2}\Big).$
Use the distribution $h_{\beta}$ to define the mixed coefficients
$\overline{I}_{l,r}^{(\beta)}:=\int_{0}^{\infty}I_{l,r}(c)\,h_{\beta}(c)\,dc,\qquad 1\leq r\leq l.$
Then one has the closed form
$\overline{I}_{l,r}^{(\beta)}=2^{-(l-r)}\,\frac{\big(\frac{\beta+1}{2}\big)_{r-1,\;1/2}}{\big(\frac{\beta}{2}+1\big)_{l-1}},$ (18)
where $(a)_{k,\,1/2}:=\prod_{i=0}^{k-1}(a+i/2)$. Consequently, the mixed EPPF
$\overline{p}(n_{1},\dots,n_{r}):=\overline{I}_{l,r}^{(\beta)}\prod_{i=1}^{r}(2n_{i}-3)!!,\qquad l=\sum_{i=1}^{r}n_{i},$
coincides with the Pitman–Yor [21] Poisson–Dirichlet EPPF $\mathrm{PD}(1/2,\beta/2)$.
Proof: Using the integral representation of $I_{l,r}(c)$ and Fubini,
$\overline{I}_{l,r}^{(\beta)}=\frac{1}{(2l-r-2)!\,Z_{\beta}}\int_{0}^{\infty}\!\!\int_{0}^{\infty}c^{\beta+r-1}\sigma^{2l-r-2}\exp\!\Big(-\frac{c^{2}+\sigma^{2}}{2}-c\sigma\Big)\,d\sigma\,dc.$
Since $\exp\!\big(-\frac{c^{2}+\sigma^{2}}{2}-c\sigma\big)=\exp\!\big(-\frac{(c+\sigma)^{2}}{2}\big)$, the change of variables $u=c+\sigma$, $w=c/(c+\sigma)$ gives a product of a Beta integral in $w$ and a Gamma integral in $u$, yielding (18). Finally, $(2n-3)!!=2^{n-1}(1/2)_{n-1}$ makes the factor $2^{-(l-r)}$ in (18) cancel, leaving exactly the standard $\mathrm{PD}(1/2,\beta/2)$ form. ∎
Excursions above 0 and the induced partition
Under the plane order induced by the leaf order as above, the contour height process $H_{\Upsilon_{L}}$ has a canonical decomposition into excursions.
Let $H:\{0,\dots,2a\}\to\operatorname{\mathbb{Z}}_{\geq 0}$ be the contour height of $\Upsilon_{L}$. Each time the traversal leaves $\Delta$ it makes a step $0\to 1$, explores one entire $\Delta$–child subtree, and then returns via a step $1\to 0$. Consequently, the number of edges from $\Delta$ equals the number of excursions of $H$ away from 0:
$r\;=\;\#\{\text{excursions of }H\text{ above }0\}.$
Each marked leaf $i\in L$ is visited during exactly one of these excursions. Thus the induced partition of $L$ (two labels are in the same block iff they lie in the same $\Delta$–component) is equivalently the partition obtained by grouping labels according to which excursion of $H_{\Upsilon_{L}}$ contains them. In particular, the limiting Gibbs($\tfrac{1}{2}$) partition obtained in this section can be viewed as the limiting excursion partition of the marked leaves for the contour process in the regime $\kappa=c\sqrt{N}$.
3.4 Trivial regimes away from the critical scale
Write $c_{N}:=\kappa_{N}/\sqrt{N}$.
Two marked vertices.
Take $l=2$. There are only two partitions: one block ($r=1$) and two singletons ($r=2$). Since $C_{2,1}=C_{2,2}=1$, Theorem 3.1 gives
$\operatorname{\mathbb{P}}(1\not\sim 2)\ \longrightarrow\ I_{2,2}(c),\qquad\operatorname{\mathbb{P}}(1\sim 2)\ \longrightarrow\ I_{2,1}(c)=1-I_{2,2}(c).$
Using the one-dimensional representation in Theorem 3.1 (valid for $l\geq 2$),
$I_{2,2}(c)=c\int_{0}^{\infty}e^{-\sigma^{2}/2-c\sigma}\,d\sigma,\qquad I_{2,1}(c)=\int_{0}^{\infty}\sigma\,e^{-\sigma^{2}/2-c\sigma}\,d\sigma.$
In particular,
$0\leq I_{2,2}(c)\leq c\int_{0}^{\infty}e^{-\sigma^{2}/2}\,d\sigma=c\sqrt{\frac{\pi}{2}},$
so $I_{2,2}(c)\to 0$ as $c\downarrow 0$. Moreover,
$0\leq I_{2,1}(c)\leq\int_{0}^{\infty}\sigma e^{-c\sigma}\,d\sigma=\frac{1}{c^{2}},$
hence $I_{2,2}(c)=1-I_{2,1}(c)\to 1$ as $c\uparrow\infty$. Therefore:
  • if $c_{N}\to 0$ (equivalently $\kappa_{N}=o(\sqrt{N})$), then $\operatorname{\mathbb{P}}(1\sim 2)\to 1$;
  • if $c_{N}\to\infty$ (equivalently $\kappa_{N}\gg\sqrt{N}$), then $\operatorname{\mathbb{P}}(1\not\sim 2)\to 1$.
One marked vertex and the distance scale.
[…]
$f_{c}(\sigma)=(\sigma+c)e^{-\sigma^{2}/2-c\sigma}\,\mathbf{1}_{\{\sigma\geq 0\}},$
Moreover, for every $t\geq 0$,
$\operatorname{\mathbb{P}}\!\left(\frac{U_{N}}{\sqrt{N}}\geq t\right)\ \longrightarrow\ \int_{t}^{\infty}f_{c}(\sigma)\,d\sigma=e^{-t^{2}/2-ct}.$
In particular, if $c_{N}\to\infty$ then for every fixed $t>0$ the right-hand side tends to 0, hence $U_{N}/\sqrt{N}\to 0$ in probability.
Conclusion.
It follows directly from these remarks that:
a) If $c_{N}\to 0$, the probability that $\Upsilon_{L}$ is a tree converges to one.
b) If $c_{N}\to\infty$, the probability that $\Upsilon_{L}$ is composed of $l$ branches intersecting only in $\Delta$ and of length $o(\sqrt{N})$ converges to one.
4 Relation to previous work
The main relations are with some results involving Brownian motion: As $l=|L|$ increases, our results in the fixed-$\kappa$ regime yield a consistent family of distributions on $\mathbb{R}_{+}$-extensions of binary tree classes with $l$ leaves. This coincides with the family of “proper $k$-trees” introduced by Aldous in [3, §4.3]. It is shown in [3] that this family admits a representation in terms of the Brownian continuum random tree (CRT). Other constructions of the CRT are given in [1, 2] using Brownian excursions or branching processes; see also [13, 14, 10].
[…]
On the complete graph, the $\kappa$-biased forest measure considered here coincides with the arboreal gas (the $q\to 0$ random-cluster model); see, for instance, Duminil-Copin, Sidoravicius and Tassion [7], and Bauerschmidt, Crawford, Helmuth and Heydenreich [4]. Related weighted forest measures were also investigated by Broder and Diaconis [5], who analyse non-uniform spanning trees and forests via determinantal formulas.
Pitman [19] introduced a coalescent construction of uniform random forests with a prescribed number of rooted trees, providing a dynamic description of forests with fixed component count. Luczak and Pittel [15] consider forests chosen uniformly among all forests with a prescribed number of components $r$, and derive detailed asymptotics for component sizes, in particular in the regime $r\asymp\sqrt{N}$. Since our $\kappa$-biased model assigns weight $\kappa^{\,r-1}$ to forests with $r$ components, it may be viewed as a mixture over $r$ of theL Luczak–Pittel uniform model, but we do not condition on the number of components and instead analyse the unconditioned model in different scaling regimes of $\kappa$.
Martin and Yeo [16] study critical random forests arising from Erdős–Rényi random graphs conditioned to be acyclic, and obtain global scaling limits for component sizes in the critical window. Pavlov [17] surveys earlier results on limit distributions for component sizes, distances and extremal tree quantities in random forests, primarily in the uniform setting.
The present paper complements the above literature by focusing on the finite-dimensional local structure of the $\kappa$-biased forest on $K_{N}$: for a fixed marked set $L$, we obtain in an entirely elementary way explicit asymptotics for the induced partition, an explicit limiting density for the rescaled branch lengths, and an asymptotic binary-tree description (with independence across blocks in the critical regime).